Suppose you’re given there is a memory access timeof 356 nanoseconds and an average page-fault servicetime of 5milliseconds.a.If one access out of 100causes a page fault, what will be the Effective Access Time? b.How does this compare to main memory access time?
Answer #1

Effective access time = (1-p)*memory access time + p*page fault time

Given memory access time = 356 nano seconds

given average page fault-service time = 5 milliseconds.

Given one access out of 100 causes a page fault i.e,

page fault rate =1/100=0.01

Effective memory access time = (1-0.01)*(356*10^-9) + (0.01)*(5*10^-2)

                                                 =0.0005003

                                                = 5.35 micro seconds

B)if the effective memory time (5.35 microseconds) compared to memory access time(356 nano seconds) then effective memory time is 1.5 times greater than memory access time